OpenStudy (anonymous):

help please Solve for V -8v+3(v-8)=21

OpenStudy (welshfella):

first distribute the 3 over the parenthes -8v +3*v + 3*-8 = 21 -8v + 3v - 24 = 21 can you continue?

OpenStudy (welshfella):

first add like terms on the left hand side

OpenStudy (anonymous):

i don't get it